2.5.


Make a prediction on what will happen if you will double (a) the length of the wire and (b) the cross-sectional area of the wire. Then, support your answer by solving for the resistance in each of the given situation.


1
Expert's answer
2021-03-17T19:47:16-0400

Resistance is given by-

"R=\dfrac{\rho l}{A}"


(i) When length become dobled than new resistance,

"R'=\dfrac{\rho \times 2l}{A}=2R"


Resistance become doubled.


(ii) When cross sectional are become doubled-


"R'=\dfrac{\rho l}{2A}=\dfrac{R}{2}"

Resistance become half.